Use a suitable identity to get each of the following products.
( 2a – 7 )(2a- 7)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

( 2a – 7 )(2a- 7)=(2a- 7)^2
= 4a^2–28a + 49
Using (a-b)^2= a^2+ b^2- 2ab

Thanks